50 ml of mineral oil is mixed with a liter of water and is subjected to high shear to create 100 nanometer spherical oil droplets. Calculate the new surface area of droplets in square meters and the number of droplets in the resulting emulsion. Assume the density of mineral oil to be 0.85 gms/cc.
Answer #1

Volume of oil should be constant

volume of oil 50 mL = n x volume of 1 drop oil

100 nm = 100 x 10-7cm = 10-5cm diameter ,

radius half of diameter = 100/2 nm = 50nm = 50 x 10-7 cm = 5 X 10-6 cm

volume of 1 drop = 4/3 x pi x r3

= 4/3 x 22/7 x (5 x 10-6)3 = 5.225 x 10-16 cm3

50 mL = 50 cm3

n = 50 cm3/(5.225 x 10-16 cm3) = 9.57 x 1016

surface area of 1 droplet = 4 x pi x r2 = 4 x 3.14 x (5 x 10-6)2 = 3.14 x 10-10 cm2

surface area of all droplets = 9.57 x 1016 x 3.14 x 10-10 = 30.05 x 106 cm2

= 3005 m2